OpenStudy (anonymous):

You currently drive 300 miles per week in a car that gets 15 miles per gallon of gas. You are considering buying a new fuel-efficient car for 12,000 (after trade-in on your current car) that gets 50 miles per gallon. Insurance premiums for the new and old car are $800 and $600 per year, respectively. You anticipate spending $1200 per year on repairs for the old car and having no repairs on the new car. Assume gas costs $3.50 per gallon. Over a five year period, is it less expensive to keep your old car or buy the new car?

OpenStudy (pfenn1):

Wow. Lots to take into account. Let's start with the old car and how much that car will cost you over a five year period. You currently drive 300 miles/week and get 15 miles/gallon and gas costs $3.50 a gallon. So over 5 years \[\text {Gas costs = }300 \frac{\text{ miles}}{\text{ week}} \times 52 \frac{\text{ weeks}}{\text{ year}} \times 5\text { years}\times \frac{\text{ gallon}}{\text{15 miles}}\times \frac{$3.50}{\text {gallon}}\]

OpenStudy (anonymous):

Thanks for posting. I believe so, I'm trying to find the chapter this equation comes from (studying for final, I haven't seen this formula in more than a few weeks now). Can you tell me the name of the formula?

OpenStudy (pfenn1):

I didn't know it had a name. I just wrote an equation that added up all the costs for the old car for 5 years.

OpenStudy (anonymous):

Well the equation you wrote looks fine. I'll continue looking for this in my textbook, I did see a question similar to this one in it. I'll bump this again if I have no luck. Thanks for the help pfenn1. I appreciate it.

OpenStudy (pfenn1):

you're welcome.

OpenStudy (anonymous):

@pfenn1

OpenStudy (pfenn1):

What's up?

OpenStudy (anonymous):

Can you take the time out to explain how to correctly solve this. I see you put the equation for the old car. Can I answer it now, or should I make the same equation for the new one first?

OpenStudy (pfenn1):

I think we need the cost for the new car first. So costs for the new car=cost of buying the new car+gas costs for 5 years +insurance costs for 5 years + 0 repair costs. Did I forget anything?

OpenStudy (anonymous):

no you got it.

OpenStudy (pfenn1):

Do you want to try to write the equation?

OpenStudy (anonymous):

Yes, though it will probably be wrong. Give me a moment to do it

OpenStudy (pfenn1):

You can do it.

OpenStudy (anonymous):

gas costs= 300x5x52x50x3.50 ? I am looking for gas costs right? Do i need to include the insurance premium and cost of the car?

OpenStudy (pfenn1):

You are close on gas costs. It is the same as I derived above except that the new car would get 50 miles to the gallon.\[\text {Gas costs (new) = }300 \frac{\text{ miles}}{\text{ week}} \times 52 \frac{\text{ weeks}}{\text{ year}} \times 5\text { years}\times \frac{\text{ gallon}}{\text{50 miles}}\times \frac{$3.50}{\text {gallon}}\]\[=\frac{(300)(52)(5)(3.50)}{50}\]

OpenStudy (pfenn1):

Btw. I just realize that I made a mistake in my equation for the cost for the old car. It should have been:\[C _{old}=\frac{(300)(52)(5)(3.50)}{15}+800(5)+1200(5)\]The first term is the gas costs, the second is the insurance for 5 years, and the last term is the repairs for 5 years.

OpenStudy (pfenn1):

So costs for new car = gas costs + insurance cost + price to buy car.

OpenStudy (anonymous):

oh...i see my mistakes. Where did you get the 52 in the gas costs though?

OpenStudy (pfenn1):

\[\text {Gas costs (new) = }300 \frac{\text{ miles}}{\text{ week}} \times 52 \frac{\text{ weeks}}{\text{ year}} \times 5\text { years}\times \frac{\text{ gallon}}{\text{50 miles}}\times \frac{$3.50}{\text {gallon}}\]

OpenStudy (anonymous):

oh lol the weeks

OpenStudy (pfenn1):

\[C _{new}=\frac{(300)(52)(5)(3.50)}{50}+600(5)+12000\]

OpenStudy (pfenn1):

So you want to calculate Cold and Cnew to see if the new or old car is less expensive.

OpenStudy (anonymous):

my first attempt at answering old car gas costs: $28200.

OpenStudy (anonymous):

is that right?

OpenStudy (pfenn1):

That's what I got.

OpenStudy (anonymous):

alright so now onto the new car...

OpenStudy (anonymous):

new car: $33200. ?

OpenStudy (anonymous):

Did i get it wrong?

OpenStudy (pfenn1):

I got $5,460 for gas costs, $3,000 for insurance costs, and $12,000 for buying the car. That doesn't add up to $33200 but I don't know what you got wrong.

OpenStudy (anonymous):

I did it right

OpenStudy (anonymous):

this time i did

OpenStudy (anonymous):

added up, the car is $20,460

OpenStudy (anonymous):

So if I did my math right, the new car is cheaper by $7,740

OpenStudy (pfenn1):

That is what I got.

OpenStudy (pfenn1):

yep

OpenStudy (anonymous):

:O I owe you, thank you for helping me with this! I can understand if you don't feel like doing another problem with me, but I have more to do. Would you mind helping me out a bit further?

OpenStudy (pfenn1):

I would like to but it is very late where I am and I have to get up early in the morning. I'm really sorry. If you need help another time, don't be afraid to ask.

OpenStudy (anonymous):

Thanks man you've done enough, have a good night!